OpenStudy (anonymous):

Use the square root 2x^(2)-14=0

OpenStudy (anonymous):

to find x?

OpenStudy (anonymous):

2x^2=14 x^2=7 x=sqrt(7) or (- sqrt(7))

OpenStudy (anonymous):

x=+-squareroot7